OpenStudy (anonymous):

Solve for the variable in the following equation. 2/3x+5/6x=1/2x+10 Choices are: x=10 or x=5/12 or x=7/12 or x=15

OpenStudy (anonymous):

2/3x can be re-written as 4/6x. 1/2x can be re-written as 3/6x. After doing this, you can begin to solve the equation more easily. 4/6x + 5/6x = 3/6x + 10. 9/6x = 3/6x + 10. 6/6x = 10. 6/6 = 1, so x = 10.

OpenStudy (anonymous):

So the final answer is 10?

OpenStudy (mertsj):

\[6\times\frac{2}{3}x+6\times\frac{5}{6}x=6\times\frac{1}{2}x+6\times10\]

OpenStudy (anonymous):

So is the answer 10 then?

OpenStudy (mertsj):

\[4x+5x=3x+60\]

OpenStudy (mertsj):

\[9x=3x+60\]

OpenStudy (mertsj):

\[6x=60\]

OpenStudy (anonymous):

So the answer is 10?

OpenStudy (mertsj):

\[x=10\]